Respected sir. When a ball is thrown up,time taken to reach the max. height and, time taken to reach the ground back will be equal, but why ?

Respected sir.
When a ball is thrown up,time taken to reach the max. height and, time taken to reach the ground back will be equal, but why ?

Grade:11

1 Answers

Khimraj
5 years ago
Lets understand mathematically
lets thrown velocity is u
then
to reach max height
0 = u -gt1
then t1 = u/g
and max height = u2/2g
for its return journy
h = u2/2g = ½*g*t22
then t2 = u/g.
so t1 = t2
Hope it clears.